the sum of Mrs.Black's age and her daughter's age is 56 years. In 8 years, Mrs. Black will be twice as old as her daughter. How old is her daughter now?
Answer by rapaljer(4671)

Let x = age of the daughter now.
y = age of Mrs. Black now.
x+8 = daughter's age in 8 years
y+8 = Mrs. Black's age in 8 years

Two equations are needed, since there are two unknowns:
Sum of the ages now is 56.
x+ y = 56

Mrs. Black will be twice as old as her daughter.
y+8 = 2(x+8) Solve for y:
y +8 = 2x + 16
y+8-8 = 2x + 16-8
y = 2x + 8

Substitute into
x+y = 56
x+(_____) = 56
x+ (2x +8)= 56
3x + 8=56
2x = 48
x= 16 Daughter's age now


x + y = 56
16 + y = 56
y = 40 = Mother's age now

In 8 years, daughter will be 24.
In 8 years, Mother will be 48.

Check is that in 8 years, mother will be twice as old as daughter.

R^2 at SCC